OpenStudy (anonymous):

find the inverse for f(x) = -3x +2 and state whether or not its a function. Can someone please help?? I have all the way through solving for x, just cant figure out how to finish it!! thanks!!

OpenStudy (anonymous):

are u asking is it inverse function?

OpenStudy (anonymous):

Find f^-1(x) for orignal equation.

OpenStudy (anonymous):

y=-3x+2 3x=2-y x=2-x/3